The length of a rectangle is 5 more than the side of a square and the width is 1 more than the side of the square. If the area of the rectangle is 96, find the length of the side of the square.
Found 2 solutions by dfrazzetto, solver91311:
Answer by dfrazzetto(283)

s = side of square
l = s + 5
w = s + 1
lxw = 96
(s + 5)(s + 1) = 96
s^2 + 6s + 5 = 96
s^2 + 6s - 91 = 0
(s + 13)(s - 7) = 0
s = -13, 7
sides of a square cannot be negative, so
s = 7

Answer by solver91311(24713)



Let represent the measure of one side of the square. Then the length and width of the rectangle must be and .

Since the area of a rectangle is given by length times width,



Multiply the binomials, put the quadratic into standard form, factor, and solve for the positive root (the negative root is absurd in the context of the measure of the side of a square).

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism